y= -3/5x
x-2y= -22
so what would be the solution to the system of equations?

Answer by Alan3354(69443)

y= -3/5x
x-2y= -22
---------------
Sub for y in the 2nd eqn.
x - 2y = -22
x -2(-3x/5) = -22
Solve for x
Then solve for y